# Style Injection in Diffusion: A Training-free Approach for Adapting Large-scale Diffusion Models for Style Transfer

### 11 Dec 2023 v1

@@ Abstract
Despite the impressive generative capabilities of diffusion models, existing diffusion model-based style transfer methods require inference-stage optimization (e.g. fine-tuning or textual inversion of style) which is time-consuming, or fails to leverage the generative ability of large-scale diffusion models. To address these issues, we introduce a novel artistic style transfer method based on a pre-trained large-scale diffusion model without any optimization. Specifically, we manipulate the features of self-attention layers as the way the cross-attention mechanism works; in the generation process, substituting the key and value of content with those of style image. This approach provides several desirable characteristics for style transfer including 1) preservation of content by transferring similar styles into similar image patches and 2) transfer of style based on similarity of local texture (e.g. edge) between content and style images. Furthermore, we introduce query preservation and attention temperature scaling to mitigate the issue of disruption of original content, and initial latent Adaptive Instance Normalization (AdaIN) to deal with the disharmonious color (failure to transfer the colors of style). Our experimental results demonstrate that our proposed method surpasses state-of-the-art methods in both conventional and diffusion-based style transfer baselines. ${}^{\ast}$ Equal contribution ${}^{\star}$ Corresponding author
@@ 2.2 Attention-based Image Editing in DM
Following the remarkable advances achieved by pre-trained text-to-image DMs [32, 42], there have been numerous image editing works [3, 19, 6, 37] utilizing these DMs. Notably, Prompt-to-Prompt [14] proposed text-based local image editing by manipulating the cross-attention map. Specifically, they observe that cross-attention largely contributes to modeling the relation between the spatial layout of the image to each word in the prompt. Hence, they substitute the original words and cross-attention map with desirable ones, obtaining edited images matched with text conditions. Subsequently, Plug-and-play [41] introduces text-guided image-to-image translation method. They found that the spatial features (i.e. feature from residual block) and self-attention map determine the spatial layout of the synthesized image. Thus, while generating a new image with the given text condition, they guide the diffusion model with features and attention map from the original image for preserving the original spatial layout.
Along with these works, we recognize the potential of attention maps in representing spatial information. However, different from the aforementioned methods concentrating on exploiting the textual condition, we focus on conditioning by style and content images. By combining the features in self-attention layers of both style and content images, we transfer the texture of the content image to the given style.
@@ 5 Experiments
[…]
ArtFID $\downarrow$ 28.801 31.420 34.685 30.720 34.605 31.282 30.350 34.630 31.856 30.933
FID $\downarrow$ 18.131 19.760 20.395 18.890 20.062 18.199 18.658 21.252 19.022 18.242
LPIPS $\downarrow$ 0.5055 0.5135 0.6212 0.5445 0.6430 0.6293 0.5439 0.5562 0.5910 0.6076
CFSD $\downarrow$ 0.2281 0.2464 0.2918 0.3011 0.3346 0.3043 0.2862 0.2920 0.3600 0.3155
Table 1: Quantitative comparison with conventional style transfer baselines.
@@ 5.2 Evaluation Protocol
Content Feature Structural Distance (CFSD)
In the style transfer evaluation, the assessment of content fidelity often relies on the LPIPS distance. However, since LPIPS utilizes the feature space of AlexNet [20] pre-trained for classification task on ImageNet [7], which is known as texture-biased [12]. Thus, the style information of the images can affect the LPIPS score. To mitigate this style influence, we additionally introduce Content Feature Structural Distance (CFSD) which is a distance measure that only considers the spatial correlation between image patches.
@@
Figure 6: Qualitative comparison with conventional baselines
@@ 5.3 Quantitative Comparison
Comparison with Conventional Style Transfer.
As shown in Tab. 1, our method largely surpasses the conventional style transfer methods in terms of ArtFID, which is known as coinciding the human preference. In addition, the proposed method records the lowest FID, which denotes that stylized images highly resemble the target styles. For content fidelity metrics, ours shows superior scores in both CFSD and LPIPS. We point out that ours achieves much lower CFSD compared to other methods, which is the metric to only consider the spatial correlation.
@@
Comparison with Diffusion-based Style Transfer.
Our method demonstrates the best performance in terms of LPIPS, FID, and their combination (ArtFID) with a large margin, as shown in Tab. 2. One significant factor for diffusion models is their running times, since they require several steps to synthesize a single image and it requires inevitable time cost. Hence, we measure the inference time for a pair of content and style images on a single TITAN RTX GPU. Our method requires a total of 12.4 seconds, with 8.2 seconds for DDIM inversions and 4.2 seconds for sampling costs. As reported, we validate the proposed method significantly faster than other methods, even exploiting large-scale DM. This faster speed comes from the fact that the proposed methods can use the much smaller steps of DDIM inversion, because we additionally utilize the features collected during inversion steps, largely reducing the necessity for perfect inversion of content and style.
Metric Time (sec) ArtFID FID LPIPS
DiffuseIT [22] 792.8 40.721 23.065 0.6921
InST [53] 816.9 40.633 21.571 0.8002
DiffStyle [18] 355.9 41.464 20.903 0.8931
Ours 12.4 28.801 18.131 0.5055
Table 2: Quantitative comparison with diffusion-based baselines
Figure 7: Qualitative comparisons with diffusion-based baselines
@@ 5.4 Qualitative Comparison
Comparison with Conventional Style Transfer.
As shown in Fig. 6, we observe that our method tends to highly preserve the structural information of the content image, while also transferring the style well. For instance, as shown in the third row, ours retains the structure of the bridge, but the baselines struggle to preserve structure or transfer the style.
Comparison with Diffusion-based Style Transfer.
We also compare our method with recent diffusion-based style transfer baselines [53, 18, 23]. As shown in Fig. 7, we observe the proposed technique transfers the style to the content well. On the other hand, baselines often lose the structure of content or fail to transfer the style, when an arbitrary content style pair is given. For instance, DiffuseIT and DiffStyle suffer from generating shape and visually plausible images or drop the original content. Differently, InST synthesizes the realistic images, while struggling to transfer style ($1^{\text{st}}$ row) or change content of image ($2^{\text{nd}},3^{\text{rd}}$ rows).
@@ 5.6 Additional Analysis
Content-Style Trade-Off.
Our proposed method offers flexible control of the trade-off relation between content and style fidelity by adjusting the parameter $\gamma$, as discussed in Sec. 4.1. In detail, we compute the FID and LPIPS while varying $\gamma$ within the range $[0.3,1]$ with a step size of 0.1. As shown in Fig. 10 (a), our method surpasses baseline methods across all ranges of content and style fidelity. This result implies the proposed method significantly outperforms the other methods, when we match style or content metric to the compared model by adjusting the $\gamma$ of ours. Note that, dotted lines refer to our model reported in Tab. 1.
@@
Study on the value of $\tau$
We observe that the gradual increase of $\tau$ enhances the performance of style transfer, although its effects on enhancement become smaller as $\tau$ goes larger, as shown in Fig. 10 (b). This result implies that the attention temperature scaling effectively works with a simple modification of the magnitude of the attention map.
Comparison with text-guided style transfer.
We additionally compare the proposed method with the style transfer methods [41, 21] which are conditioned on the textual inputs. As text-guided methods tend to modify the style largely, we use $\gamma=0.3$ for this experiment. Since the text condition hardly contains all the information in the style image such as texture and color tones, the transferred results less resemble the target style, as shown in Fig. 11. Differently, we validate that the proposed method successfully transfers the style with high fidelity.
@@ 6 Conclusion
Appendices
Ablation study for color transfer capability.
To validate the efficacy of the ablated methods for color transfer, we employ the RGB-$uv$ histogram proposed in HistoGAN [1] to measure color transfer capability. Specifically, for a given input image $I$, we convert it into the log-chroma space. For example, choosing the R color channel as the primary and normalizing by G and B yields:
@@
Qualitative comparison with ablation of attention temperature scaling.
To highlight the effects of attention temperature scaling, we provide some examples of style transfer results while ablating the attention scaling. As shown in Fig. 12, we validate that the attention scaling makes the model to synthesize sharp images and well-preserve the patterns in the given style image (e.g. stars in left example). This experimental result confirms the significance of the proposed attention temperature scaling method. Note that, we use $\gamma=0.3$ for this experiment, to keep the strong effect of style transfer in visualization.
Additional qualitative results.
We additionally compare the proposed method with the most recent baseline (AesPA-Net) and baseline with the lowest ArtFID (AdaAttN). As shown in Fig. 1314, we observe that ours better-transfers the local texture of a given style into the content image.

### 20 Mar 2024 v2

@@ Abstract
Despite the impressive generative capabilities of diffusion models, existing diffusion model-based style transfer methods require inference-stage optimization (e.g. fine-tuning or textual inversion of style) which is time-consuming, or fails to leverage the generative ability of large-scale diffusion models. To address these issues, we introduce a novel artistic style transfer method based on a pre-trained large-scale diffusion model without any optimization. Specifically, we manipulate the features of self-attention layers as the way the cross-attention mechanism works; in the generation process, substituting the key and value of content with those of style image. This approach provides several desirable characteristics for style transfer including 1) preservation of content by transferring similar styles into similar image patches and 2) transfer of style based on similarity of local texture (e.g. edge) between content and style images. Furthermore, we introduce query preservation and attention temperature scaling to mitigate the issue of disruption of original content, and initial latent Adaptive Instance Normalization (AdaIN) to deal with the disharmonious color (failure to transfer the colors of style). Our experimental results demonstrate that our proposed method surpasses state-of-the-art methods in both conventional and diffusion-based style transfer baselines. Codes are available at github.com/jiwoogit/StyleID.
${}^{\ast}$ Equal contribution ${}^{\dagger}$ Corresponding author
@@ 2.2 Attention-based Image Editing in DM
Following the remarkable advances achieved by pre-trained text-to-image DMs [35, 45], there have been numerous image editing works [3, 20, 7, 40] utilizing these DMs. Notably, Prompt-to-Prompt [15] proposed text-based local image editing by manipulating the cross-attention map. Specifically, they observe that cross-attention largely contributes to modeling the relation between the spatial layout of the image to each word in the prompt. Hence, they substitute the original words and cross-attention map with desirable ones, obtaining edited images matched with text conditions. Subsequently, Plug-and-play [44] introduces text-guided image-to-image translation method. They found that the spatial features (i.e. feature from residual block) and self-attention map determine the spatial layout of the synthesized image. Thus, while generating a new image with the given text condition, they guide the diffusion model with features and attention map from the original image for preserving the original spatial layout. Recently, MasaCtrl [4] proposes mutual self-attention control for consistent image editing using text prompts. In detail, they retain the source image’s $key$ and $value$ of the self-attention layers, while conditioning the model with desired text prompts.
Along with these works, we recognize the potential of attention maps in representing spatial information. However, different from the aforementioned methods concentrating on exploiting textual condition, we focus on conditioning by style and content images composed of two images from distinct styles. By combining the features in self-attention layers of both style and content images with precise adjustment of statistics in intermediate representations, we transfer the texture of the content image to the given style.
@@ 4.2 Attention Temperature Scaling
[…]
ArtFID $\downarrow$ 28.801 31.420 34.685 30.720 34.605 31.282 30.350 34.630 31.856 30.933 40.721 40.633 41.464
FID $\downarrow$ 18.131 19.760 20.395 18.890 20.062 18.199 18.658 21.252 19.022 18.242 23.065 21.571 20.903
LPIPS $\downarrow$ 0.5055 0.5135 0.6212 0.5445 0.6430 0.6293 0.5439 0.5562 0.5910 0.6076 0.6921 0.8002 0.8931
CFSD $\downarrow$ 0.2281 0.2464 0.2918 0.3011 0.3346 0.3043 0.2862 0.2920 0.3600 0.3155 0.3428 0.6759 0.2819
[…]
@@ 5.2 Evaluation Protocol
Content Feature Structural Distance (CFSD). In the style transfer evaluation, the assessment of content fidelity often relies on the LPIPS distance. However, since LPIPS utilizes the feature space of AlexNet [21] pre-trained for classification task on ImageNet [8], which is known as texture-biased [13]. Thus, the style information of the images can affect the LPIPS score. To mitigate this style influence, we additionally introduce Content Feature Structural Distance (CFSD) which is a distance measure that only considers the spatial correlation between image patches.
@@
[…]
Figure 7: Qualitative comparison with best ArtFID (AdaAttN) and most recently proposed baselines (AesPA-Net) with additional zooming details
@@ 5.3 Quantitative Comparison
Comparison with Conventional Style Transfer. As shown in Tab. 1, our method largely surpasses the conventional style transfer methods in terms of ArtFID, which is known as coinciding the human preference. In addition, the proposed method records the lowest FID, which denotes that stylized images highly resemble the target styles. For content fidelity metrics, ours shows superior scores in both CFSD and LPIPS. We point out that ours achieves much lower CFSD compared to other methods, which is the metric to only consider the spatial correlation.
@@
Comparison with Diffusion-based Style Transfer. Our method demonstrates the best performance in terms of LPIPS, FID, and their combination (ArtFID) with a large margin, as shown in Tab. 1. One significant factor for diffusion models is their running times, since they require several steps to synthesize a single image and it requires inevitable time cost. Hence, we measure the inference time for a pair of content and style images on a single TITAN RTX GPU, as shown in Tab. 2. Our method requires a total of 12.4 seconds, with 8.2 seconds for DDIM inversions and 4.2 seconds for sampling costs. As reported, we validate the proposed method significantly faster than other methods, even exploiting large-scale DM. This faster speed comes from the fact that the proposed methods can use the much smaller steps of DDIM inversion, because we additionally utilize the features collected during inversion steps, largely reducing the necessity for perfect inversion of content and style.
Metric DiffuseIT InST DiffStyle Ours
Time (sec) 792.8 816.9 355.9 12.4
Table 2: Comparison of inference time of diffusion-based methods for style transferring a given style and content pair
@@ 5.4 Qualitative Comparison
Comparison with Conventional Style Transfer. As shown in Fig. 6, we observe that our method tends to highly preserve the structural information of the content image, while also transferring the style well. For instance, as shown in the third row, ours retains the structure of the bridge, but the baselines struggle to preserve structure or transfer the style. We also provide the qualitative comparison with zooming details in Fig. 7 and Supplementary.
Comparison with Diffusion-based Style Transfer. We also compare our method with recent diffusion-based style transfer baselines [56, 19, 24]. As shown in Fig. 6, we observe the proposed technique transfers the style to the content well. On the other hand, baselines often lose the structure of content or fail to transfer the style, when an arbitrary content style pair is given. For instance, DiffuseIT and DiffStyle suffer from generating shape and visually plausible images or drop the original content. Differently, InST synthesizes the realistic images, while struggling to transfer style ($1^{\text{st}}$ row) or change content of image ($2^{\text{nd}},3^{\text{rd}}$ rows).
@@ 5.6 Additional Analysis
Content-Style Trade-Off. Our proposed method offers flexible control of the trade-off relation between content and style fidelity by adjusting the parameter $\gamma$, as discussed in Sec. 4.1. In detail, we compute the FID and LPIPS while varying $\gamma$ within the range $[0.3,1]$ with a step size of 0.1. As shown in Fig. 10 (a), our method surpasses baseline methods across all ranges of content and style fidelity. This result implies the proposed method significantly outperforms the other methods, when we match style or content metric to the compared model by adjusting the $\gamma$ of ours. Note that, dotted lines refer to our model reported in Tab. 1.
@@
Study on the value of $\tau$. We observe that the gradual increase of $\tau$ enhances the performance of style transfer, although its effects on enhancement become smaller as $\tau$ goes larger, as shown in Fig. 10 (b). This result implies that the attention temperature scaling effectively works with a simple modification of the magnitude of the attention map.
Comparison with text-guided style transfer. We additionally compare the proposed method with the style transfer methods [44, 22] which are conditioned on the textual inputs. As text-guided methods tend to modify the style largely, we use $\gamma=0.3$ for this experiment. Since the text condition hardly contains all the information in the style image such as texture and color tones, the transferred results less resemble the target style, as shown in Fig. 11. Differently, we validate that the proposed method successfully transfers the style with high fidelity.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by MSIT/IITP (No. 2022-0-00680, 2019-0-00421, 2020-0-01821, 2021-0-02068), and MSIT&KNPA/KIPoT (Police Lab 2.0, No. 210121M06).
Supplementary Material
7 Appendix
Ablation study for color transfer capability. To validate the efficacy of the ablated methods for color transfer, we employ the RGB-$uv$ histogram proposed in HistoGAN [1] to measure color transfer capability. Specifically, for a given input image $I$, we convert it into the log-chroma space. For example, choosing the R color channel as the primary and normalizing by G and B yields:
@@ 7 Appendix
Qualitative comparison with ablation of attention temperature scaling. To highlight the effects of attention temperature scaling, we provide some examples of style transfer results while ablating the attention scaling. As shown in Fig. 12, we validate that the attention scaling makes the model to synthesize sharp images and well-preserve the patterns in the given style image (e.g. stars in left example). This experimental result confirms the significance of the proposed attention temperature scaling method. Note that, we use $\gamma=0.3$ for this experiment, to keep the strong effect of style transfer in visualization.
Quantitative comparison in the other set. In Tab. 5, we conduct quantitative experiments on a new set of style-content pairs (20 contents, 40 styles) randomly sampled without any overlap with original images. As reported, the performance enhancement of the proposed method still holds, confirming hyperparameters are well-generalized.
LPIPS is affected by texture and color, as it is based on CNN features [13]. To evalute the content and color independently, we measure LPIPS-Grayscale and Histogram-loss in supplementary against the recent and lowest ArtFID baselines (AesPA-Net, InST, AdaAttN). As reported in Tab. 5, ours achieves lowest LPIPS-Grayscale, and highest color similarity.
Ours AesPA-Net AdaAttN InST
ArtFID 30.38 34.55 31.87 39.11
FID 18.87 21.09 19.21 20.46
LPIPS 0.528 0.563 0.576 0.822
LPIPS-Gray 0.417 0.443 0.450 0.731
Histogram-loss 0.303 0.321 0.331 0.653
Table 5: Quantitative comparison in newly sampled test set.
Analysis on feature space of query preservation. Fig. 13 visualizes features of $Q^{c}_{t}$, $Q^{s}_{t}$, $Q^{cs}_{t}$, and $\tilde{Q}^{cs}_{t}$ for a style-content pair. As shown, interpolated features ($\tilde{Q}^{cs}_{t}$) are located in in-distribution nearby contents, since we gradually combine content query ($Q^{c}_{t}$) and stylized one ($Q^{cs}_{t}$) along with entire reverse process.
Furthermore, we compute the average distance of $\tilde{Q}^{cs}_{t}$ toward top-5 Nearest Neighbors (NNs) in (content, style, itself) and the number of them in NNs for all injected layers with $t$=[10, 20, 30, 40]. Distances and # NNs are (5.49, 9.06, 4.43), (1.24, 0.00, 3.76), implying $\tilde{Q}^{cs}_{t}$ residing in in-distribution nearby content.
[…]
Style transfer with text prompts. In this paragraph, We exploit text prompt, obtained by BLIP [26], for DDIM inversion instead of null text token. Images in ‘data_vis’ in the official repository are used, in which easy to caption as they mostly consist of single object. As a result, ours w/ text shows slight improvement as in Tab. 6.
User study. We compare ours with AesPA-Net and InST, the most recent conventional and diffusion methods, for 18 users and 10 examples per user. We observe that (57.2%, 76.7%) of users prefer the proposed method over (AesPA-Net, InST). Note that, ours has a much faster inference speed than InST.
ArtFID FID LPIPS-Gray
Ours (w/ empty prompt, default) 34.9 21.2 0.47
Ours (w/ BLIP prompt) 34.5 20.9 0.47
Table 6: Ablation study of the null text token in the diffusion process.
Qualitative comparison with StyleDiffusion. As the implementation of StyleDiffusion [48] is unavailable, we compare ours with examples in supplementary of StyleDiffusion [48]. We obtain style-content pairs of StyleDiffusion in repositories of their baselines. We observe that ours is more suitable for transferring local textures, while StyleDiffusion tends to change the structure of the image significantly, as shown in Fig. 15. We hypothesize that optimizing the style in CLIP [34]’s semantically rich feature space forces StyleDiffusion to be trained in that manner.
Figure 14: Qualitative comparisons with diffusion-based baselines
Figure 15: Qualitative comparison with StyleDiffusion. * denotes cropped version of images. We use $\gamma=0.5$ for visualization.
Additional qualitative results. We additionally compare the proposed method with the most recent baseline (AesPA-Net) and baseline with the lowest ArtFID (AdaAttN). Fig. 14 shows the additional qualitative comparison of ours with diffusion model baselines. Moreover, as shown in Fig. 1617, we observe that ours better-transfers the local texture of a given style into the content image.